OpenStudy (anonymous):

Homework11 and LAB11answers......!

OpenStudy (anonymous):

p1: L=0.65mH, RL=4 ohms, RC=490Kohms 1.)1/(RC*C) + RL/L 2.)92.236 (f=650khz) 3.)4.5 4.)22 (f=1330khz) 5.)16 p2: 1.)vm*0.1=0.01 2.)Cs/Rp=2.55 3.)Cs/Rp=2.55 4.)0.8 p3: 1.)C 2.)R 3.)L 4.)4 5.)L 6.)1000 7.)1 8.)16

OpenStudy (anonymous):

LAB11: 1.)0.001592 2.)1.592e-7 3.)0.0356 4.)10

OpenStudy (anonymous):

hi rvamsi, what is the eq. for H11P2 part (4) ...

OpenStudy (anonymous):

no wonder why i was having trouble with the ifrst part of lab 11. Those guys are idiots. They said wo=10k which means 2*pi*f=10k.. they actually meant that fresonance=10k, THEREFORE wo=62.8k

OpenStudy (anonymous):

Q1 and Q2 equations please my frequency is different

OpenStudy (anonymous):

hi vamsi can u please say me the equation for finding capacitor in H11Q1

OpenStudy (anonymous):

for H11P1 que1 use C=1/(L*((2*3.14*f)^2))

OpenStudy (anonymous):

Hello there fellas, can't get H11P2 - part 4 and 5. can anybody sort this for me? I have f=1150 kHz. Appreciate it

OpenStudy (anonymous):

problem 2 or problem 1?

OpenStudy (anonymous):

Oops, sorry, P1 it is. part 4 and 5

OpenStudy (anonymous):

At f = 1150 kHz, finding capacitance and badwidth

OpenStudy (anonymous):

ans for 4=29.496 ans for 5=75.343

OpenStudy (anonymous):

please select best response if it is correct if not pls respond i will try again

OpenStudy (anonymous):

Ahh, thank you Mr. Siva, part 4 is correct but nit agreeing on 5. the one that asks for bandwidth. Do you know what could be wrong? Thank you sir

OpenStudy (anonymous):

oops small mistake ans for 5 is 11.991

OpenStudy (anonymous):

Grrreeeat! Thank you very much. You have helped me ace it. :)

OpenStudy (anonymous):

No mention if u are done with H12 P3 please give me the formula or equations for that

OpenStudy (anonymous):

Plz can you tell me the formula to calculate bandwidth for part 3 of HW 11

OpenStudy (anonymous):

and I cant understand how to find Cp..

OpenStudy (anonymous):

hey what is the need of BW in P3

OpenStudy (anonymous):

pls give me the Q no

OpenStudy (anonymous):

BW for part 3 of H11P1...but I got it...I wasn't dividing the answer with 2*pie so....but now I am stuck in H11P2 last part

OpenStudy (anonymous):

P1 2&3 f= 1010 4&5 1030 khz

OpenStudy (anonymous):

b) (thats the formula = 1/(2*pi*f)^(2)*L) c) (thats the formula = 1/(RC*C) + RL/L) d) (thats the formula = 1/(2*pi*f)^(2)*L) e) c) (thats the formula = 1/(RC*C) + RL/L) H11P2 d) 0.1*VM(t)

OpenStudy (anonymous):

Click best if helped I got 38.2 and for b and d of h11p2. Cant get right answer for c and e Plz Help Rc = 490k, RL = 4 L = .65mH f= 1010 and 1030 rsp.

OpenStudy (anonymous):

Coolampio, I try to use the formula 1/(RC*C) + RL/L) in c) and e), but I've got a wrong answer. my f=740kHz & 1660kHz. I've got capacitances 71pF & 14pF respectively. But in questions c & e I've got 35kHz and 151kHz, it's wrong.

OpenStudy (anonymous):

For d, I got 36.7 forgot to mention up having pblem with same ones my self

OpenStudy (anonymous):

I find it! For c) & e) parts of h11p1 we need divide the formula: "1/(RC*C) + RL/L" by 2pi.

OpenStudy (anonymous):

What is the bandwidth, in 770 kHz, of the tank at kHz?

OpenStudy (anonymous):

Mahmoudmosad, your bandwidth is 5.9kHz